# Is Chain-of-Thought Reasoning of LLMs a Mirage? A Data Distribution Lens

### v3

@@ 5 Task Generalization
Proposition 5.1 (Task Generalization Complexity).
@@
$\operatorname{TGC}(C)=$ $\alpha\sum_{i=1}^{m}\mathbb{I}\left[a_{i}\notin\mathcal{E}^{i}_{\text{train }}\right]+\beta\sum_{j=1}^{n}\mathbb{I}\left[f_{j}\notin\mathcal{F}_{\text{train }}\right]+\gamma\mathbb{I}\left[\left(f_{1},f_{2},\ldots,f_{k}\right)\notin\mathcal{P}_{\text{train }}\right]+C_{T}$ (11)
@@
Theorem 5.1 (Task Generalization Failure Threshold).
@@
$P(\text{correct}|C)\leq e^{-\delta(\text{TGC}(C)-\tau)}$ (12)
The proof is provided in Appendix A.2.
5.1 Transformation Generalization
Transformation generalization evaluates the ability of CoT reasoning to effectively transfer when models encounter novel transformations during testing, which is an especially prevalent scenario in real-world applications.
Figure 3: Performance of CoT reasoning on transformation generalization. Efficacy of CoT reasoning declines as the degree of distributional discrepancy increases.
Experimental Setup. To systematically evaluate the impact of transformations, we conduct experiments by varying transformations between training and testing sets while keeping other factors constant (e.g., elements, length, and format). Guided by the intuition formalized in Proposition 5.1, we define four incremental levels of distribution shift in transformations as shown in Figure 2: (i) In-Distribution (ID): The transformations in the test set are identical to those observed during training, e.g., $f_{1}\circ f_{1}\rightarrow f_{1}\circ f_{1}$. (ii) Composition (CMP): Test samples comprise novel compositions of previously encountered transformations, though each individual transformation remains familiar, e.g., ${f_{1}\circ f_{1},f_{1}\circ f_{2},f_{2}\circ f_{1}}\rightarrow f_{2}\circ f_{2}$. (iii) Partial Out-of-Distribution (POOD): Test data include compositions involving at least one novel transformation not seen during training, e.g., $f_{1}\circ f_{1}\rightarrow f_{1}\circ f_{2}$. (iv) Out-of-Distribution (OOD): The test set contains entirely novel transformation types that are unseen in training, e.g., $f_{1}\circ f_{1}\rightarrow f_{2}\circ f_{2}$.
Table 1: Full chain evaluation under different scenarios for transformation generalization.
@@ 5.1 Transformation Generalization
Transformation (Train $\rightarrow$ Test) Scenario Exact Match Edit Distance BLEU Score
$f_{1}\circ f_{1}\rightarrow f_{1}\circ f_{1}$ ID 100.00% 0 1
$\{f_{2}\circ f_{2},f_{1}\circ f_{2},f_{2}\circ f_{1}\}\rightarrow f_{1}\circ f_{1}$ CMP 0.01% 0.1326 0.6867
$f_{1}\circ f_{2}\rightarrow f_{1}\circ f_{1}$ POOD 0.00% 0.1671 0.4538
$f_{2}\circ f_{2}\rightarrow f_{1}\circ f_{1}$ OOD 0.00% 0.2997 0.2947
@@
Findings. Figure 3 illustrates the performance of the full chain under different distribution discrepancies computed by task generalize complexities (normalized between 0 and 1) in Definition 5.1. We can observe that, in general, the effectiveness of CoT reasoning decreases when distribution discrepancy increases. For the instance shown in Table 1, from in-distribution to composition, POOD, and OOD, the exact match decreases from 1 to 0.01, 0, and 0, and the edit distance increases from 0 to 0.13, 0.17 when tested on data with transformation $f_{1}\circ f_{1}$. Apart from ID, LLMs cannot produce a correct full chain in most cases, while they can produce correct CoT reasoning when exposed to some composition and POOD conditions by accident. As shown in Table 2, from $f_{1}\circ f_{2}$ to $f_{2}\circ f_{2}$, the LLMs can correctly answer 0.1% of questions. A close examination reveals that it is a coincidence, e.g., the query element is A, N, A, N, which happened to produce the same result for the two operations detailed in the Appendix D.1. When further analysis is performed by breaking the full chain into reasoning steps and answers, we observe strong consistency between the reasoning steps and answers. For example, under the composition generalization setting, the reasoning steps are entirely correct on test data distribution $f_{1}\circ f_{1}$ and $f_{2}\circ f_{2}$, but with wrong answers. Probe these insistent cases in Appendix D.1, we can find that when a novel transformation (say $f_{1}\circ f_{1}$) is present, LLMs try to generalize the reasoning paths based on the most similar ones (i.e., $f_{1}\circ f_{2}$) seen during training, which leads to correct reasoning paths, yet incorrect answer, which echo the example in the introduction. Similarly, generalization from $f_{1}\circ f_{2}$ to $f_{2}\circ f_{1}$ or vice versa allows LLMs to produce correct answers that are attributed to the commutative property between the two orthogonal transformations with unfaithful reasoning paths. Collectively, the above results indicate that the CoT reasoning fails to generalize to novel transformations, not even to novel composition transforms. Rather than demonstrating a true understanding of text, CoT reasoning under task transformations appears to reflect a replication of patterns learned during training.
Table 2: Evaluation on different components in CoT reasoning on transformation generalization. CoT reasoning shows inconsistency with the reasoning steps and answers.
Transformation (Train $\rightarrow$ Test) Exact Match Edit Distance
Reason Answer Full Chain Reason Answer Full Chain
$\{f_{1}\circ f_{1},f_{1}\circ f_{2},f_{2}\circ f_{1}\}\rightarrow f_{2}\circ f_{2}$ 100.00% 0.01% 0.01% 0.000 0.481 0.133
$\{f_{1}\circ f_{2},f_{2}\circ f_{1},f_{2}\circ f_{2}\}\rightarrow f_{1}\circ f_{1}$ 100.00% 0.01% 0.01% 0.000 0.481 0.133
$f_{1}\circ f_{2}\rightarrow f_{2}\circ f_{1}$ 0.00% 100.00% 0.00% 0.373 0.000 0.167
$f_{2}\circ f_{1}\rightarrow f_{1}\circ f_{2}$ 0.00% 100.00% 0.00% 0.373 0.000 0.167
Figure 4: Performance on unseen transformation using SFT in various levels of distribution shift. Introducing a small amount of unseen data helps CoT reasoning to generalize across different scenarios.
Experiment settings. To further probe when CoT reasoning can generalize to unseen transformations, we conduct supervised fine-tuning (SFT) on a small portion $\lambda$ of unseen data. In this way, we can decrease the distribution discrepancy between the training and test sets, which might help LLMs to generalize to test queries.
Findings. As shown in Figure 4, we can find that generally a very small portion ($\lambda=1.5e^{-4}$) of data can make the model quickly generalize to unseen transformations. The less discrepancy between the training and testing data, the quicker the model can generalize. This indicates that a similar pattern appears in the training data, helping LLMs to generalize to the test dataset.
5.2 Element Generalization
@@
Findings. Similar to transformation generalization, the performances degrade sharply when facing the distribution shift consistently across all transformations, as shown in Figure 5. From ID to CMP and OOD, the exact match decreases from 1.0 to 0 and 0, for all cases. Most strikingly, the BLEU score is 0 when transferred to $f_{1}$ and $f_{2}$ transformations. A failure case in Appendix D.1 shows that the models cannot respond to any words when novel elements are present. We further explore when CoT reasoning can generalize to novel elements by conducting SFT. The results are summarized in Figure 6. We evaluate the performance under three exact matches for the full chain under three scenarios, CMP based on the edit distance n. The result is similar to SFT on transformation. The performance increases rapidly when presented with similar (a small $n$) examples in the training data. Interestingly, the exact match rate for CoT reasoning aligns with the lower bound of performance when $n=3$, which might suggest the generalization of CoT reasoning on novel elements is very limited, even SFT on the downstream task. When we further analyze the exact match of reasoning, answer, and token during the training for $n=3$, as summarized in Figure 6b. We find that there is a mismatch of accuracy between the answer and the reasoning step during the training process, which somehow might provide an explanation regarding why CoT reasoning is inconsistent in some cases.
@@
6 Length Generalization
Length Extrapolation Bound. We establish a power-law relationship for length extrapolation:
Proposition 6.1 (Length Extrapolation Gaussian Degradation).
The proof is provided in Appendix A.3.
Table 3: Evaluation for text length generalization.
Length Exact Match (%) Edit Distance BLEU Score
Full Chain Reason Answer Full Chain Reason Answer Full Chain Reason Answer
2 0.00% 0.00% 0.00% 0.3772 0.4969 0.5000 0.4214 0.1186 0.0000
3 0.00% 0.00% 0.00% 0.2221 0.3203 0.2540 0.5471 0.1519 0.0000
4 100.00% 100.00% 100.00% 0.0000 0.0000 0.0000 1.0000 1.0000 1.0000
5 0.00% 0.00% 0.00% 0.1818 0.2667 0.2000 0.6220 0.1958 0.2688
6 0.00% 0.00% 0.00% 0.3294 0.4816 0.3337 0.4763 0.1174 0.2077
6.1 Text Length Generalization
@@
Findings. As illustrated in the Table 3, the CoT reasoning failed to directly generate two test cases even though those lengths present a mild distribution shift. Further, the performance declines as the length discrepancy increases shown in Figure 7. For instance, from data with $l=4$ to those with $l=3$ or $l=5$, the BLEU score decreases from 1 to 0.55 and 0.62. Examples in Appendix D.1 indicate that LLMs attempt to produce CoT reasoning with the same length as the training data by adding or removing tokens in the reasoning chains. The efficacy of CoT reasoning length generalization deteriorates as the discrepancy increases. Moreover, we consider using a different padding strategy to decrease the divergence between the training data and test cases. We found that padding to the max length doesn’t contribute to length generalization. However, the performance increases when we replace the padding with text by using the group strategy, which indicates its effectiveness.
6.2 Reasoning Step Generalization
The reasoning step generalization investigates whether models can extrapolate to reasoning chains requiring different steps $k$ from those observed during training. which is a popular setting in multi-step reasoning tasks.
Experiment settings. Similar to text length generalization, we first pre-train the LLM with reasoning step $k=2$, and evaluate on data with reasoning step $k=1$ or $k=3$.
(a) Reasoning step. From k=2 to k=1
(b) Reasoning step. From k=2 to k=3
Figure 8: Test performance for reasoning-step generalization across varying training data compositions. Performance varies systematically with changes in the distribution of training data components.
Findings. As showcased in Figure 8, CoT reasoning cannot generalize across data requiring different reasoning steps, indicating the failure of generalization. Then, we try to decrease the distribution discrepancy introduced by gradually increasing the ratio of unseen data while keeping the dataset size the same when pre-training the model. And then, we evaluate the performance on two datasets. As we can observe, the performance on the target dataset increases along with the ratio. At the same time, the LLMs can not generalize to the original training dataset because of the small amount of training data. The trend is similar when testing different-step generalization, which follows the intuition and validates our hypothesis directly.
7 Format Generalization
Definition 7.1 (Format Alignment Score).
For training prompt distribution $P_{train}$ and test prompt $p_{test}$:
where $\phi$ is a prompt embedding function.
(a) Format generalization. Performance under various perturbation methods.
(b) Format generalization. Performance vs. various applied perturbation areas.
Figure 9: Performance of format generalization. Testing performance varies with different noise levels and areas where the noise is applied.
Experiment settings. To systematically probe this, we introduce four distinct perturbation modes to simulate scenario in real-world: (i) insertion, where a noise token is inserted before each original token; (ii) deletion: it deletes the original token; (iii) modification: it replaces the original token with a noise token; and (iv) hybrid mode: it combines multiple perturbations. Each mode is applied for tokens with probabilities $p$, enabling us to quantify the model’s resilience to increasing degrees of prompt distribution shift.
Findings. As shown in Figure 9a, we found that generally CoT reasoning can be easily affected by the format changes. No matter insertion, deletion, modifications, or hybrid mode, it creates a format discrepancy that affects the correctness. Among them, the deletion slightly affects the performance. While the insertions are relatively highly influential on the results. We further divide the query into several sections: elements, transformations, and prompt tokens. As shown in Figure 9b, we found that the elements and transformation play an important role in the format, whereas the changes to other tokens rarely affect the results.
8 Temperature and Model Size
@@
9 Discussion and Implication
Prioritize Out-of-Distribution (OOD) Testing. Standard validation practices, where the test set closely mirrors the training set, are insufficient to gauge the true robustness of a CoT-enabled system. Practitioners must implement rigorous adversarial and OOD testing that systematically probes for vulnerabilities across task, length, and format variations.
10 Conclusion
In this paper, we critically examine the COT reasoning of LLMs through the lens of data distribution, revealing that the perceived structured reasoning capability largely arises from inductive biases shaped by in-distribution training data. We propose a controlled environment, DataAlchemy, allowing systematic probing of CoT reasoning along three crucial dimensions: task structure, reasoning length, and query format. Empirical findings consistently demonstrate that CoT reasoning effectively reproduces reasoning patterns closely aligned with training distributions but suffers significant degradation when faced with distributional deviations. Such observations reveal the inherent brittleness and superficiality of current CoT reasoning capabilities. We provide insights that emphasize real-world implications for both practitioners and researchers.
[…]
Appendix A Proof of Theorems
A.1 Proof of CoT Generalization Bound

### v4

@@ E.1 Task Distribution
Proposition E.1 (Task Generalization Complexity).
@@
$\operatorname{TGC}(C)=\ \alpha\sum_{i=1}^{m}\mathbb{1}\left[a_{i}\notin\mathcal{E}^{i}_{\text{train }}\right]+\beta\sum_{j=1}^{n}\mathbb{1}\left[f_{j}\notin\mathcal{F}_{\text{train }}\right]+\gamma\mathbb{1}\left[\left(f_{1},f_{2},\ldots,f_{k}\right)\notin\mathcal{P}_{\text{train }}\right]+C_{T}$ (14)
@@
Theorem E.1 (Task Generalization Failure Threshold).
@@
$P(C)\leq e^{-\delta(\text{TGC}(C)-\tau)}$ (15)
The proof is provided in Appendix H.2.
E.2 Length Distribution
Proposition E.2 (Length Extrapolation Gaussian Degradation).
The proof is provided in Appendix H.3.
E.3 Format Distribution
Definition E.1 (Format Alignment Score).
For training prompt distribution $P_{train}$ and test prompt $p_{test}$:
where $\phi$ is a prompt embedding function.
Table 4: Evaluation on transformation generalization.
@@ E.3 Format Distribution
Transformation (Train $\rightarrow$ Test) Exact Match (%) Edit Distance BLEU Score
Reasoning Answer Full Chain Full Chain Reasoning Answer Full Chain Reasoning Answer
$\{f_{2}\circ f_{3},\;f_{3}\circ f_{2},\;f_{3}\circ f_{3}\}\rightarrow f_{2}\circ f_{2}$ 6.66 10.25 6.66 0.0941 0.0718 0.2244 0.5417 0.6683 0.1982
$\{f_{2}\circ f_{3},\;f_{3}\circ f_{2},\;f_{2}\circ f_{2}\}\rightarrow f_{3}\circ f_{3}$ 100.00 9.19 9.19 0.0488 0.0000 0.1768 0.8220 1.0000 0.1932
$f_{2}\circ f_{3}\rightarrow f_{3}\circ f_{2}$ 0.00 0.00 0.00 0.2997 0.3728 0.4808 0.2000 0.0019 0.0000
$f_{3}\circ f_{2}\rightarrow f_{2}\circ f_{3}$ 0.00 0.00 0.00 0.2334 0.2249 0.4808 0.2548 0.0952 0.0000
Appendix F Additional Quantitative Results
F.1 Transformation Generalization
F.1.1 Detailed analysis.
@@ F.1 Transformation Generalization
For the instance shown in Table 1, from in-distribution to composition, POOD, and OOD, the exact match decreases from 1 to 0.01, 0, and 0, and the edit distance increases from 0 to 0.13, 0.17 when tested on data with transformation $f_{1}\circ f_{1}$. Apart from ID, LLMs cannot produce a correct full chain in most cases, while they can produce correct CoT reasoning when exposed to some composition and POOD conditions by accident. As shown in Table 2, from $f_{1}\circ f_{2}$ to $f_{2}\circ f_{2}$, the LLMs can correctly answer 0.1% of questions. A close examination reveals that it is a coincidence, e.g., the query element is A, N, A, N, which happened to produce the same result for the two operations detailed in the Appendix G. When further analysis is performed by breaking the full chain into reasoning steps and answers, we observe strong consistency between the reasoning steps and answers. For example, under the composition generalization setting, the reasoning steps are entirely correct on test data distribution $f_{1}\circ f_{1}$ and $f_{2}\circ f_{2}$, but with wrong answers. Probe these insistent cases in Appendix G, we can find that when a novel transformation (say $f_{1}\circ f_{1}$) is present, LLMs try to generalize the reasoning paths based on the most similar ones (i.e., $f_{1}\circ f_{2}$) seen during training, which leads to correct reasoning paths, yet incorrect answer, which echo the example in the introduction. Similarly, generalization from $f_{1}\circ f_{2}$ to $f_{2}\circ f_{1}$ or vice versa allows LLMs to produce correct answers that are attributed to the commutative property between the two orthogonal transformations with unfaithful reasoning paths. Collectively, the above results indicate that the CoT reasoning fails to generalize to novel transformations, not even to novel composition transforms. Rather than demonstrating a true understanding of text, CoT reasoning under task transformations appears to reflect a replication of patterns learned during training.
F.1.2 Introducing another transformation.
Our main conclusion is that CoT reasoning cannot generalize to genuinely novel transformations, including unseen compositions, even when the underlying primitives are well learned. Counterexamples based on commutativity have already substantiated this claim (Table 2). We now extend the analysis by introducing a non-commutative transformation $f_{3}$ and evaluating generalization behaviors that cannot be trivially explained by commutative equivalence. The results are summarized in Table 4.
When models are trained on mixtures of transformations involving $f_{2}$ and $f_{3}$ and evaluated on unseen compositions (e.g., ${f_{2}\circ f_{3},f_{3}\circ f_{2},f_{3}\circ f_{3}}\rightarrow f_{2}\circ f_{2}$), performance remains extremely poor across all metrics. Exact match accuracy for the full chain stays below 10%, and both edit distance and BLEU scores indicate substantial divergence from the correct reasoning traces and final answers. Notably, even when the reasoning component occasionally achieves high exact match (e.g., 100% reasoning accuracy in ${f_{2}\circ f_{3},f_{3}\circ f_{2},f_{2}\circ f_{2}}\rightarrow f_{3}\circ f_{3}$), the corresponding full-chain and answer-level accuracy collapse, revealing a clear disconnect between locally plausible reasoning steps and globally correct execution.
More strikingly, in strictly non-commutative transfer settings such as $f_{2}\circ f_{3}\rightarrow f_{3}\circ f_{2}$ and its reverse, the model fails completely: exact match drops to 0 across reasoning, answer, and full chain, while edit distance sharply increases and BLEU scores approach zero. Unlike earlier commutativity-induced cases—where incorrect reasoning paths could still yield correct answers—these failures demonstrate that once superficial equivalences are removed, CoT reasoning no longer exhibits any meaningful transfer. This provides strong evidence that prior apparent generalization was not driven by learning transformation semantics, but rather by exploiting distributional artifacts such as commutativity.
Overall, Table 4 reinforces our central claim: CoT reasoning does not support systematic generalization to novel transformations. Instead, its success hinges on structural overlap and distributional shortcuts present in the training data. When these shortcuts are eliminated via non-commutative transformations, both reasoning traces and answers degrade simultaneously, exposing the brittleness of CoT reasoning under genuine task-level distribution shifts. Illustrative examples of $f_{3}$ are provided in Appendix B for completeness
F.2 Element Generalization
@@
Findings. Similar to transformation generalization, the performances degrade sharply when facing the distribution shift consistently across all transformations, as shown in Figure 11. From ID to CMP and OOD, the exact match decreases from 1.0 to 0 and 0, for all cases. Most strikingly, the BLEU score is 0 when transferred to $f_{1}$ and $f_{2}$ transformations. A failure case in Appendix G shows that the models cannot respond to any words when novel elements are present. We further explore when CoT reasoning can generalize to novel elements by conducting SFT. The results are summarized in Figure 12. We evaluate the performance under three exact matches for the full chain under three scenarios, CMP based on the edit distance n. The result is similar to SFT on transformation. The performance increases rapidly when presented with similar (a small $n$) examples in the training data. Interestingly, the exact match rate for CoT reasoning aligns with the lower bound of performance when $n=3$, which might suggest the generalization of CoT reasoning on novel elements is very limited, even SFT on the downstream task. When we further analyze the exact match of reasoning, answer, and token during the training for $n=3$, as summarized in Figure 12b. We find that there is a mismatch of accuracy between the answer and the reasoning step during the training process, which somehow might provide an explanation regarding why CoT reasoning is inconsistent in some cases.
Table 5: Evaluation on text length generalization.
Text Length Exact Match (%) Edit Distance BLEU Score
Reasoning Answer Full Chain Full Chain Reasoning Answer Full Chain Reasoning Answer
2 0.00 0.00 0.00 0.3772 0.4969 0.5000 0.4214 0.1186 0.0000
3 0.00 0.00 0.00 0.2221 0.3203 0.2540 0.5471 0.1519 0.0000
4 100.00 100.00 100.00 0.0000 0.0000 0.0000 1.0000 1.0000 1.0000
5 0.00 0.00 0.00 0.1818 0.2667 0.2000 0.6220 0.1958 0.2688
6 0.00 0.00 0.00 0.3294 0.4816 0.3337 0.4763 0.1174 0.2077
F.3 Text Length Generalization
@@
Findings. As illustrated in the Table 5, the CoT reasoning failed to directly generate two test cases even though those lengths present a mild distribution shift. Further, the performance declines as the length discrepancy increases shown in Figure 13. For instance, from data with $l=4$ to those with $l=3$ or $l=5$, the BLEU score decreases from 1 to 0.55 and 0.62. Examples in Appendix G indicate that LLMs attempt to produce CoT reasoning with the same length as the training data by adding or removing tokens in the reasoning chains. The efficacy of CoT reasoning length generalization deteriorates as the discrepancy increases. Moreover, we consider using a different padding strategy to decrease the divergence between the training data and test cases. We found that padding to the max length doesn’t contribute to length generalization. However, the performance increases when we replace the padding with text by using the group strategy, which indicates its effectiveness.
F.4 Temperature and Model Size
@@
Figure 15: Task, length, and format generalization of LLMs with settings. The data distribution lens is invariant across LLMs with various sizes and architectures.
Table 6: Task generalization performance of SOTA LLMs (mean $\pm$ std).
Model Scenario Exact Match (%) Edit Distance BLEU Score
LLaMA3-8B ID $100.00\pm 0.00$ $0.00\pm 0.00$ $1.00\pm 0.00$
CMP $8.52\pm 0.00$ $0.23\pm 0.00$ $0.61\pm 0.00$
POOD $0.00\pm 0.01$ $0.25\pm 0.01$ $0.46\pm 0.00$
OOD $0.00\pm 0.00$ $0.27\pm 0.01$ $0.27\pm 0.00$
Qwen3-14B-Instruct ID $100.00\pm 0.00$ $0.00\pm 0.00$ $1.00\pm 0.00$
CMP $0.01\pm 0.01$ $0.17\pm 0.02$ $0.61\pm 0.00$
POOD $0.00\pm 0.00$ $0.26\pm 0.01$ $0.42\pm 0.00$
OOD $0.00\pm 0.00$ $0.38\pm 0.01$ $0.36\pm 0.00$
@@
Table 7: Reasoning step generalization of SOTA LLMs (mean $\pm$ std).
Model Reasoning Step ($k$) Exact Match (%) Edit Distance BLEU Score
LLaMA3-8B 1 $0.00\pm 0.00$ $0.75\pm 0.01$ $0.18\pm 0.00$
2 $100.00\pm 0.00$ $0.00\pm 0.00$ $1.00\pm 0.00$
3 $0.00\pm 0.00$ $0.54\pm 0.01$ $0.40\pm 0.00$
Qwen3-14B-Instruct 1 $0.00\pm 0.00$ $0.54\pm 0.02$ $0.35\pm 0.00$
2 $100.00\pm 0.00$ $0.00\pm 0.00$ $1.00\pm 0.00$
3 $0.26\pm 0.05$ $0.65\pm 0.08$ $0.20\pm 0.00$
Table 8: Format generalization under different noise levels for SOTA LLMs (mean $\pm$ std).
Model Noise Level (%) Exact Match (%) Edit Distance BLEU Score
LLaMA3-8B 0 $100.00\pm 0.00$ $0.00\pm 0.00$ $1.00\pm 0.00$
5 $32.30\pm 0.47$ $41\pm 0.02 0.0$$04\pm 0.00$
10 $17.74\pm 0.38$ $45\pm 0.00 0.0$$04\pm 0.00$
15 $9.27\pm 0.29$ $49\pm 0.05 0.0$$04\pm 0.00$
20 $4.72\pm 0.21$ $53\pm 0.03 0.0$$03\pm 0.00$
25 $2.32\pm 0.15$ $57\pm 0.00 0.0$$03\pm 0.00$
30 $1.09\pm 0.10$ $60\pm 0.08 0.0$$03\pm 0.00$
Qwen3-14B-Instruct 0 $100.00\pm 0.00$ $0.00\pm 0.00$ $1.00\pm 0.00$
5 $9.96\pm 0.30$ $0.22\pm 0.03$ $0.41\pm 0.00$
10 $5.09\pm 0.22$ $0.35\pm 0.01$ $0.24\pm 0.00$
15 $2.43\pm 0.15$ $0.44\pm 0.02$ $0.16\pm 0.00$
20 $1.19\pm 0.11$ $0.52\pm 0.05$ $0.11\pm 0.00$
25 $0.53\pm 0.07$ $57\pm 0.04 0.0$$08\pm 0.00$
30 $0.22\pm 0.05$ $62\pm 0.02 0.0$$06\pm 0.00$
F.5 Internal Validity
Figures 8 and 15 illustrate task, length, and format generalization across a wide range of GPT- and LLaMA-style models. Increasing distribution discrepancy leads to a monotonic degradation of CoT performance regardless of architectural choice or parameter count. While larger models achieve uniformly higher absolute scores under near in-distribution conditions, they do not exhibit qualitatively different robustness profiles under moderate or severe shifts. This suggests that the observed failures of CoT reasoning cannot be attributed to insufficient capacity or architectural idiosyncrasies, but rather reflect a shared inductive bias learned from training distributions.
Model scaling consistently improves performance in ID and mildly shifted regimes (e.g., CMP or small reasoning-step extrapolation), but provides diminishing returns as the discrepancy increases toward POOD and OOD settings. In particular, larger models tend to preserve fluent intermediate reasoning traces even when final answers deteriorate, mirroring the same failure modes observed in smaller models. This pattern reinforces the interpretation that scaling primarily enhances pattern interpolation within the support of the training distribution, rather than enabling principled extrapolation beyond it.
Notably, the qualitative nature of errors remains stable across model sizes and architectures. Under task shifts, models consistently default to the closest seen transformation pattern; under length shifts, they bias toward producing training-length reasoning chains; and under format perturbations, they remain sensitive to surface-level noise in structurally salient regions. The persistence of these behaviors across settings further supports the internal validity of the data distribution lens: CoT reasoning behaves as a distribution-sensitive generative process rather than an architecture-specific reasoning mechanism.
Taken together, these results demonstrate that our core findings are not artifacts of a particular model family, scale, or training instability. Instead, the dependence of CoT effectiveness on task, length, and format distributions emerges as a stable and reproducible phenomenon across controlled LLM instantiations. This strengthens the claim that distribution discrepancy—rather than model choice—is the dominant factor governing when CoT reasoning succeeds or fails.
F.6 External Validity
As illustrated in Table 6, 7 & 8, the results on LLaMA3-8B-Instruct and Qwen3-14B-Instruct confirm that the behaviors identified in DataAlchemy persist in real-world, pretrained LLMs despite unknown and opaque training distributions. In the in-distribution (ID) setting, both models achieve perfect performance (100% exact match, zero edit distance, BLEU score of 1), indicating that the curated tasks are well within the expressive and optimization capacity of modern instruction-tuned models.
However, once distribution shifts are introduced, performance degrades sharply and systematically. Under composition (CMP), exact match accuracy drops to 8.52% for LLaMA3-8B and near zero (0.01%) for Qwen3-14B, while POOD and OOD settings result in complete failure (0% exact match) for both models. Correspondingly, edit distance increases and BLEU score decreases monotonically from CMP to OOD. These trends mirror those observed in models trained from scratch, reinforcing that the failure of CoT reasoning under task-level distribution shifts is not an artifact of synthetic training or limited model scale.
Reasoning-step generalization further highlights the limited extrapolation ability of CoT reasoning in SOTA models. Both LLaMA3-8B and Qwen3-14B achieve perfect performance at the in-distribution reasoning depth ($k=2$), but fail almost entirely at unseen depths. For $k=1$ and $k=3$, exact match accuracy collapses to 0% in nearly all cases, accompanied by large edit distances and sharply reduced BLEU scores. Notably, Qwen3-14B exhibits a marginal non-zero accuracy (0.26%) at $k=3$, but this gain is unstable and negligible relative to the ID performance. This pattern indicates that even large, instruction-tuned models do not acquire a length-agnostic or algorithmic reasoning procedure, but instead internalize a narrowly scoped reasoning template tied to the training distribution.
Format generalization experiments show consistent degradation trends across noise levels for both models. As noise increases from 0% to 30%, exact match accuracy decays smoothly from 100% to near zero, while edit distance increases and BLEU score decreases monotonically. Although LLaMA3-8B initially appears more sensitive at low noise levels (e.g., 32.30% exact match at 5% noise versus 9.96% for Qwen3-14B), both models converge to similarly poor performance under higher noise. The near-parallel degradation curves suggest that robustness to surface-level perturbations is constrained by distributional alignment.
Across all settings, the reported standard deviations are small relative to the absolute performance gaps between ID and shifted distributions. This indicates that the observed trends are stable across runs and not driven by sampling noise or stochastic decoding effects. In particular, zero or near-zero variance in ID settings confirms deterministic mastery of in-distribution patterns, while low variance under OOD conditions reflects consistently poor generalization rather than brittle or erratic behavior.
Overall, these results establish strong external validity for the proposed data distribution lens. Despite their scale, architectural sophistication, and instruction tuning, SOTA LLMs exhibit the same qualitative behaviors as controlled models trained in DataAlchemy: strong in-distribution performance, sharp degradation under task, length, and format shifts, and limited benefit from increased model capacity. This alignment between synthetic and real-world settings supports the claim that the effectiveness and brittleness of CoT reasoning are fundamentally governed by distributional alignment, rather than by pretraining scale or architectural advances alone.
Appendix G Additional Qualitative Analysis
G.1 Failures in Task Generalization
G.1.1 Failures caused by Transformation Generalization
The following case illustrates the failure caused by the transformation partial OOD setting discussed in Section C.1.1. Specifically, the model is trained exclusively on the compositional transformation $f_{1}\circ f_{2}$, but is evaluated on the unseen composition $f_{1}\circ f_{1}$.
Query:
H U S P [F1] [F1] <think>
Response:
U H F C [F2] <answer> H F C U
Ground-Truth:
U H F C [F1] <answer> H U S P
The results demonstrate that the model persists in applying the training composition ($f_{1}\circ f_{2}$) rather than adapting to the new sequence ($f_{1}\circ f_{1}$), even though the constituent transformation $f_{1}$ was observed during training. The model incorrectly generates the token [F2] in the reasoning step, mirroring the fixed pattern learned from the training data.
G.1.2 Failures caused by Element Generalization
The following case illustrates the failure caused by the element-level OOD setting discussed in Section C.1.2. Specifically, the training data was restricted to visible elements A–M, while the model was evaluated on queries consisting exclusively of elements N–Z.
Query:
N N N Q [F2] [F1] <think>
Response:
A A C E [F1] <answer> N N P R
Ground-Truth:
N N Q N [F1] <answer> A A D A
In the generated reasoning, the model incorrectly outputs AACE instead of the expected intermediate result NNQN. This error arises because all training queries consisted solely of elements A–M. Consequently, the model fails to generalize the transformation $f_{2}$ to the unseen elements; instead, it reverts to the training distribution, attempting to replicate the A–M patterns observed during training.
G.2 Failures in Length Generalization
G.2.1 Failures in Text Length Generalization
The following failure case demonstrates the model’s inability to generalize to unseen text lengths. In this experiment, the model was trained on sequences of length four but evaluated on sequences of length five.
Query:
I G L L Q [F1] [F2] <think>
Response:
T Y Y [F2] <answer> T Y Y V
Ground-Truth:
V T Y Y D [F2] <answer> T Y Y D V
Despite the increased length of the input query (five atoms), the model fails to adapt. As seen in the generated answer (T Y Y V), the model rigidly adheres to the length constraint observed during training, outputting a sequence of four atoms instead of the required five.
G.2.2 Failures in Reasoning Step Generalization
Query:
A A B D [F1] <answer>
Response:
N O A Z N N O Q [F1] <answer> A A B D
Ground-Truth:
N N O Q
G.3 Failures in Format Generalization
The following failure case demonstrates the model’s inability to generalize under format change. Specifically, here we use delete as the format-changing mechanism.
Query:
A A A T [F1] <answer>
Response:
N G N G Y
Ground-Truth:
N N N G
Appendix H Proof of Theorems
H.1 Proof of CoT Generalization Bound